# Measuring eccentricity and gas-induced perturbation from gravitational waves of LISA massive black hole binaries

### v1

@@ 3.3 Effective dephasing due to gas
Both phasing terms from migration and accretion have the same $v^{-8}$ frequency dependence with respect to the vacuum GW phase, which implies that gas corrections to the SPA phase enter at the $-4$PN order. The amplitudes ($C_{\rm mig}$ and $C_{\rm acc}$) of both effects have similar dependencies on the binary-disc parameters. However, migration dephasing becomes stronger than accretion for unequal mass ratios.
The overall phasing contribution from gas can be expressed as
$\Delta\psi_{\rm gas}=10^{-16}C_{\rm g}\psi^{(0)}_{\rm TF2}v^{-2n_{\rm g}},$ (11)
@@
$C_{\rm g}$ $=\frac{\xi}{1.0}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.25}\right)^{-2}\frac{M_{z}}{10^{5}{\rm M}_{\sun}},$ (12)
@@
where we model $C_{\rm g}$ based on the migration dephasing in Eq. (10) and any uncertainties about the dependence on the binary-disc parameters are folded into $\xi$, which can either be positive or negative.
@@ 4 Parameter space, LISA response, and time delay interferometry
To study the evolution of a MBHB embedded in an accretion disc, we mainly consider the binary-disc parameters summarized in Table 1, which are defined in the LISA frame. There are eight intrinsic parameters (first eight rows of Table 1) and six extrinsic parameters (last six rows of Table 1). We further divide the intrinsic parameters into five intrinsic-merger $\{M_{z},q,\chi_{1},\chi_{2},t_{c}\}$ and three intrinsic-inspiral $\{e_{0},C_{\rm g},n_{\rm g}\}$ parameters, due to their relative importance in different phases of the GW source evolution.
@@
$M_{z}$
Total redshifted mass
${\rm M}_{\sun}$
$q$
Mass ratio
Dimensionless
$\chi_{1,2}$
Spin parameters of both BHs
Dimensionless
$t_{c}$
Time of coalescence
year
$e_{0}$
Initial eccentricity
Dimensionless
$C_{\rm g}$
Environmental amplitude
Dimensionless
$n_{\rm g}$
Environmental semi-major axis power-law relative to GWs
Dimensionless
$D_{\rm L}$
Luminosity distance
Mpc
$\phi_{c}$
Phase at coalescence
Radian
$\imath$
Inclination
Radian
$\lambda$
Ecliptic latitude
Radian
$\beta$
Ecliptic longitude
Radian
$\psi$
Initial polarization angle
Radian
Table 1: Parameters of interest in the LISA frame.
@@
The extrinsic fiducial parameters are $z=1$, which corresponds to $D_{\rm L}=6791.3$ Mpc for the best-fit Planck Collaboration et al. (2020) cosmology, and all angles are set to $0.5$ radians. The systems we consider here spend at least 4 years in the LISA band before merging. We limit $e_{0}\geq 10^{-3}$, since eccentricities below that are unmeasurable for MBHBs (Garg et al., 2024). Considering that BHs in gas are expected to be highly spinning (see, e.g. Reynolds, 2021), we chose $\chi_{1,2}=0.9$ as our fiducial case. However, we find that our constraints on intrinsic-inspiral parameters are almost independent of the exact spin magnitude. The SNRs of the sources in the grid specified in (4) range from $\sim 150$ to $\sim 2500$ at $z=1$ with tiny differences for the two times of coalescence. [^9] See Fig. 2 of Garg et al. (2024) for SNRs measured in a one year observation for our systems of interest at different redshifts. In the next section, we compute relative errors on the measurement of parameters.
@@ 5 Fisher formalism
Using the Fisher formalism (Vallisneri, 2008), we compute the expected error ($\sigma^{\rm Fisher}_{\theta}$) on each parameter $\theta$ and set a threshold on the relative error $\delta\theta[\%]\equiv 100*\sigma^{\rm Fisher}_{\theta}/|\theta|<50$ to define when a parameter is considered to be measured. We use $50$ per cent as a threshold since we are interested in determining whether the parameter differs from zero, i.e., whether an effect is present, rather than wanting an accurate measurement of a large effect. We first consider a MBHB with $M_{z}=10^{5}~{}{\rm M}_{\sun}$ and $q=8.0$ and with a coalescence time of $t_{c}=4$ years, which has an SNR of $377.85$ and generates $19159$ GW cycles in the LISA band. We study both circular and eccentric (with $e_{0}=0.01$ or $e_{0}=0.1$) cases. In this section, we always keep the intrinsic-merger parameters $\{M_{z},q,\chi_{1},\chi_{2},t_{c}\}$ free and we show relative errors in vacuum in Table 2. Unsurprisingly, including eccentricity as a free parameter increases the uncertainties on all the intrinsic-merger parameters. Furthermore, given that we expect our fiducial CBD system to have the environmental power-law $n_{\rm g}=4$ and $n_{\rm g}$ has a non-Gaussian posterior, based on Fig. 8, we always keep it fixed to its fiducial value.
@@ 5.1.1 Circular MBHB
We compute relative errors in the presence of an environment for a circular system. In Fig. 1, we show results for $\delta C_{\rm g}[\%]$ as a function of $C_{\rm g}$. We find that $\delta C_{\rm g}[\%]\approx 5.34\times 10^{2}/|C_{\rm g}|$, almost independently of the sign of the environmental amplitude. The absolute error $\sigma^{\rm Fisher}_{C_{\rm g}}$ is independent of the value of $C_{\rm g}$, as expected due to the phase correction from the gas being linearly dependent on $C_{\rm g}$. Moreover, we can well-measure $|C_{\rm g}|\gtrsim 10$.
Figure 1: The relative error on the measurement of the environmental effective amplitude $\delta C_{\rm g}[\%]$ as a function of $C_{\rm g}$, denoted by a red ‘+’ symbol, assuming zero eccentricity and fixed environmental power-law $n_{\rm g}$. The dashed line represents the circular case. The $C_{\rm g}=0$ case is computed using $C_{\rm g}=10^{-3}$ for numerical reasons. We denote the $50$ per cent well-measured threshold by a solid gray line.
@@ 5.1.2 Eccentric MBHB
In the presence of an environment, we compute relative uncertainties for injected signals with eccentricities $e_{0}=0.01$ and $e_{0}=0.1$. In Fig. 2, we show results for $\delta C_{\rm g}[\%]$ and $\delta e_{0}[\%]$ as a function of $C_{\rm g}$. We find that $\delta C_{\rm g}[\%]\approx 6.35\times 10^{3}/|C_{\rm g}|$, irrespective of the eccentricity value, as expected, but higher than the circular case shown in Fig. 1. The eccentricity uncertainty, $\delta e_{0}[\%]$, is nearly independent of the environmental amplitude, but reaches higher values than for the vacuum case shown in Table. 2. Both increases can be attributed to having an extra free intrinsic-inspiral parameter. $|C_{\rm g}|\gtrsim 10^{2}$ is well-measured for our fiducial system.
@@ 5.1.3 Summary
  • For all the scenarios explored here, amplitudes less than $C_{\rm g}]\lesssim 10$ are not well measured.
In the next section, we will extend the Fisher results to cover the full parameter space defined in (4). Based on the results in this section, we will only consider $C_{\rm g}\geq 1$ and choose $C_{\rm g}=100$ and $e_{0}=0.1$ as our fiducial intrinsic-inspiral parameters.
@@ 5.2.1 $\delta C_{\rm g}[\%]$
We want to study the uncertainty in the environmental amplitude $C_{\rm g}$ as a function of $C_{\rm g}$ for both circular and eccentric systems. Based on the findings of Section 5.1, we only need to consider positive values of $C_{\rm g}\in\{10^{3},10^{2},10^{1},10^{0}\}$ and choose one eccentricity (here $e_{0}=0.1$). In Fig. 3, we show $\delta C_{\rm g}[\%]$ for various different choices of the intrinsic-merger parameters and strength of the environmental effect. Since $M_{z}$ defines if the inspiral part of the signal is in the low or high sensitivity region of the LISA frequency band, it affects the uncertainty on $C_{\rm g}$. Moreover, $q$ and $t_{c}$ set the number of GW cycle and so they also affect the measurement of the environmental amplitude. Based on the parameter dependence in Eq. (12), for $M_{z}=10^{6}{\rm M}_{\sun}$, $q=8.0$ or $\eta=0.1$, and $\xi=10$, we can confidently measure ${\rm f}_{\rm Edd}\gtrsim 0.1$ for circular systems and ${\rm f}_{\rm Edd}\gtrsim 1$ for eccentric MBHBs. These constraints on ${\rm f}_{\rm Edd}$ depends upon setting $\xi=10$, which is consistent with an extrapolation of the simulation results of Dittmann & Ryan (2022), who found increasingly strong torques on binaries the more thin the disc is in their Fig. 3.
Figure 3: The relative Fisher errors on the environmental amplitude, $\delta C_{\rm g}[\%]$, as a function of $|C_{\rm g}|$ and the total mass $M_{z}$. The environmental power-law $n_{\rm g}$ is fixed to $4$ in all panels. In all eight panels, we vary $M_{z}$ from $10^{4.5}$ to $10^{6}~{}{\rm M}_{\sun}$ and $C_{\rm g}$ from $10^{3}$ to $1$. For all the left panels, we set $t_{c}=1$ year and assume $t_{c}=4$ years for all the right panels. We have either $q=1.2$ (in the first and third rows) or $q=8.0$ (in the second and fourth rows). In the top four panels, we assume circular orbits and in the bottom four panels we set $e_{0}=0.1$, and allow the initial eccentricity $e_{0}$ to also be a free parameter. We draw a solid blue line to identify the region of well-measured parameters which have relative errors below $50$ per cent. Moreover, we suppress all errors beyond $100$ per cent. We can at best well measure $C_{\rm g}\gtrsim 10$ for circular binaries with $q=8$ and $t_{c}=4$ years and are not able to constrain $C_{\rm g}$ at all for eccentric systems with $q=1.2$ and $t_{c}=1$ year.
@@ 5.2.2 $\delta e_{0}[\%]$
Now we want to compute the relative errors on the initial eccentricity, $e_{0}$, as a function of $e_{0}$, with or without the environmental effect. In Fig. 4, we show $\delta e_{0}[\%]$ in vacuum and in the presence of an environment with $C_{\rm g}=100$. The measurement on $e_{0}$ strongly depends upon $M_{z}$ for a similar reason to the dependence of $\delta C_{\rm g}[\%]$ upon $M_{z}$. It also varies weakly with $q$ and $t_{c}$, since limited information is contained in the GW cycles observed from the very early inspiral.
Figure 4: The relative error on the initial eccentricity, $\delta e_{0}[\%]$, as a function of $e_{0}$ and $M_{z}$. In all eight panels we vary eccentricities from $10^{-3}$ to $0.1$ and masses from $10^{4.5}$ to $10^{6}~{}{\rm M}_{\sun}$. In the top four panels, we assume vacuum and in the bottom four panels we have $C_{\rm g}=100$ and fixed $n_{\rm g}=4$. All other parameter choices are the same as in Fig. 3. In vacuum, we are able to measure at minimum $e_{0}\gtrsim 10^{-2.75}$. While in the presence of an environment, at best we can measure $e_{0}\gtrsim 10^{-2}$.
For $t_{c}=1$ year we find a higher minimum measurable eccentricity in vacuum than what is reported in Garg et al. (2024) because of the larger number of free intrinsic-merger parameters $\{\chi_{1,2},t_{c}\}$ in the analysis.
@@ 6 Bayesian inference
We use the exact same Bayesian setup outlined in Section 4 of Garg et al. (2024) to find posterior distributions for a few systems of interest. Salient features include a zero noise realization and the Fisher initialization101010starting walkers inside a multivariate Gaussian with its mean given by the injected values and covariance by the Fisher matrix. to accelerate the parallel tempering Markov-chain Monte Carlo (PTMCMC) runs using the ptmcmc [^11] https://github.com/JohnGBaker/ptmcmc sampler. We set uniform priors for all parameters. The resulting posteriors can be used to cross-verify the Fisher matrix results reported in Section 5.
First, we show posteriors for our fiducial system in Fig. 5. We can infer that all the parameters are well-recovered and the posteriors almost overlap with the Fisher matrix predictions. We conclude that our Fisher results in Section 5 are robust. [^12] The posterior for $t_{c}=1$ year is similarly Gaussian as in Fig. 5, although with higher covariances as expected.
@@
Figure 5: Posterior distributions (solid black) for a gas-embedded eccentric MBHB with parameters $M_{z}=10^{5}~{}{\rm M}_{\sun}$, $q=8.0$, $\chi_{1,2}=0.9$, and $t_{c}=4$ years with effective gas-amplitude $C_{\rm g}=100$ and fixed power-law $n_{\rm g}=4$. The blue lines mark the injected values, the middle dashed line shows the median of the distribution, and the two extreme vertical dashed lines indicate the $90$ per cent symmetric credible interval. The two-dimensional contours of the posteriors indicate $68$, $95$, and $99$ per cent credible intervals. We also show the Fisher matrix predictions (dashed red) for comparison.
Next, we compare posteriors on the environmental amplitude for a circular system to those for two eccentric systems in Fig. 6. The $C_{\rm g}$ posteriors in all cases peak around the injected value of $100$. Also, as expected from Fig. 2, the shape of the posteriors for both non-zero eccentricities are almost the same. In the circular case, we have no support for $C_{\rm g}=0$. The broadening of the posterior is dominated by the extra degree of freedom in the model, rather than by the the presence of eccentricity in the signal itself.
@@
Last, we compare $e_{0}$ and $C_{\rm g}$ posteriors between fixed intrinsic-merger parameters, free intrinsic-merger parameters, and free intrinsic-merger and extrinsic parameters in Fig. 7.131313See Fig. E2 of Garg et al. (2024) for posteriors of the extrinsic parameters. We find that all posteriors peak around the injected values with smallest uncertainties for the fixed intrinsic-merger variables and largest uncertainties for the free extrinsic variables. These results can be attributed to having fewer or more free parameters. For fixed intrinsic-merger parameters, we have no support for $C_{\rm g}=0$. Also, the errors on the intrinsic parameters are almost independent of the inclusion or not of extrinsic parameters in the model.

### v2

@@ 3.3 Effective dephasing due to gas
Both phasing terms from migration and accretion have the same $v^{-8}$ frequency dependence with respect to the vacuum GW phase, which implies that gas corrections to the SPA phase enter at the $-4$PN order. The amplitudes ($C_{\rm mig}$ and $C_{\rm acc}$) of both effects have similar dependencies on the binary-disc parameters. However, $C_{\rm mig}/C_{\rm acc}\sim(0.25/\eta)$ implies that dephasing due to migration is only comparable to accretion dephasing for equal-mass and becomes increasingly weaker for higher mass ratio binaries.
The overall phasing contribution from gas from both migration and accretion can be expressed as
$\Delta\psi_{\rm gas}=C_{\rm g}\psi^{(0)}_{\rm TF2}v^{-2n_{\rm g}}\left(\frac{\eta}{0.1}\right)^{-2}\frac{M_{z}}{10^{5}{\rm M}_{\sun}},$ (17)
@@
$C_{\rm g}$ $=10^{-15}\frac{\xi}{1.0}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon},$ (18)
@@
We model $\Delta\psi_{\rm gas}$ based on the migration dephasing in Eq. (13) and any uncertainties about the dependence on the binary-disc parameters are folded into $\xi$, which can either be positive or negative.
@@ 4 Parameter space, LISA response, and time delay interferometry
To study the evolution of a MBHB embedded in an accretion disc, we mainly consider the binary-disc parameters summarized in Table 1, which are defined in the detector frame. There are eight intrinsic parameters (first eight rows of Table 1) and six extrinsic parameters (last six rows of Table 1). We further divide the intrinsic parameters into five intrinsic-merger $\{M_{z},q,\chi_{1},\chi_{2},t_{c}\}$ and three intrinsic-inspiral $\{e_{0},C_{\rm g},n_{\rm g}\}$ parameters, due to their relative importance in different phases of the GW source evolution.
@@
$M_{z}$ Total redshifted mass ${\rm M}_{\sun}$
$q$ Mass ratio Dimensionless
$\chi_{1,2}$ Spin parameters of both BHs Dimensionless
$t_{c}$ Time of coalescence year
$e_{0}$ Initial eccentricity Dimensionless
$C_{\rm g}$ Environmental amplitude Dimensionless
$n_{\rm g}$ Environmental semi-major axis power-law relative to GWs Dimensionless
$D_{\rm L}$ Luminosity distance Mpc
$\phi_{c}$ Phase at coalescence Radian
$\imath$ Inclination Radian
$\lambda$ Ecliptic latitude Radian
$\beta$ Ecliptic longitude Radian
$\psi$ Initial polarization angle Radian
Table 1: Parameters of interest in the detector frame.
@@
The extrinsic fiducial parameters are $z=1$, which corresponds to $D_{\rm L}=6791.3$ Mpc for the best-fit Planck Collaboration et al. (2020) cosmology, and all angles are set to $0.5$ radians. The systems we consider here spend at least 4 years in the LISA band before merging. We limit $e_{0}\geq 10^{-3}$, since eccentricities below that are unmeasurable for MBHBs (Garg et al., 2024b). Considering that BHs in gas are expected to be highly spinning (see, e.g. Reynolds, 2021), we chose $\chi_{1,2}=0.9$ as our fiducial case. However, we find that our constraints on intrinsic-inspiral parameters are almost independent of the exact spin magnitude. The SNRs of the sources in the grid specified in (4) range from $\sim 150$ to $\sim 2500$ at $z=1$ with tiny differences for the two times of coalescence. [^9] See Fig. 2 of Garg et al. (2024b) for SNRs measured in a one year observation for our systems of interest at different redshifts. In the next section, we compute relative errors on the measurement of parameters.
@@ 5 Fisher formalism
Using the Fisher formalism (Vallisneri, 2008), we compute the expected error ($\sigma^{\rm Fisher}_{\theta}$) on each parameter $\theta$ and set a threshold on the relative error $\delta\theta[\%]\equiv 100*\sigma^{\rm Fisher}_{\theta}/|\theta|<50$ to define when a parameter is considered to be measured. We use $50$ per cent as a threshold since we are interested in determining whether the parameter differs from zero, i.e., whether an effect is present, rather than wanting an accurate measurement of a large effect. We can use the Savage-Dickey ratio (Dickey, 1971) to convert our Fisher-based error threshold to the more customary Bayes factor ($\mathcal{B}$). For a recovered value $k\sigma$ away from zero assuming a Gaussian posterior with standard deviation $\sigma$ and a uniform prior of width $l\sigma$ on the parameter of interest, implies a Bayes factor of (Dickey, 1971; Garg et al., 2024b)
$\mathcal{B}=\frac{1}{l\sigma}\sqrt{2\pi\sigma^{2}}\exp\left({\frac{1}{2}\frac{(k\sigma)^{2}}{\sigma^{2}}}\right)=\frac{\sqrt{2\pi}}{l}\exp\left(\frac{k^{2}}{2}\right).$ (21)
For a $50$ per cent threshold, i.e., $k=2$ and assuming $l=5$, which ensures the prior is just big enough to contain the entire region of high likelihood support and hence is the largest Bayes factor that would not be prior-dominated, gives $\ln\mathcal{B}\approx 1.3$, which has a substantial strength (Taylor, 2021). To get decisive strength ($\ln\mathcal{B}\gtrsim 5$), we would need a relative error below $30$ percent or $k\gtrsim 3.3$.
We first consider a MBHB with $M_{z}=10^{5}~{}{\rm M}_{\sun}$ and $q=8.0$ and with a coalescence time of $t_{c}=4$ years, which has an SNR of $377.85$ and generates $19159$ GW cycles in the LISA band. We study both circular and eccentric (with $e_{0}=0.01$ or $e_{0}=0.1$) cases. In this section, we always keep the intrinsic-merger parameters $\{M_{z},q,\chi_{1},\chi_{2},t_{c}\}$ free and we show relative errors in vacuum in Table 2. Unsurprisingly, including eccentricity as a free parameter increases the uncertainties on all the intrinsic-merger parameters. Furthermore, given that we expect our fiducial CBD system to have the environmental power-law $n_{\rm g}=4$ and $n_{\rm g}$ has a non-Gaussian posterior, based on Fig. 6 discussed in Section 6, we always keep it fixed to its fiducial value.
@@ 5.1.1 Circular MBHB
We compute relative errors in the presence of an environment for a circular system. In Fig. 1, we show results for $\delta C_{\rm g}[\%]$ as a function of $C_{\rm g}$. We find that $\delta C_{\rm g}[\%]\approx 5.34\times 10^{2}/|10^{15}C_{\rm g}|$, almost independently of the sign of the environmental amplitude. The absolute error $\sigma^{\rm Fisher}_{C_{\rm g}}$ is independent of the value of $C_{\rm g}$, as expected due to the phase correction from the gas being linearly dependent on $C_{\rm g}$. Moreover, we can well-measure $|C_{\rm g}|\gtrsim 10^{-14}$.
Figure 1: The relative error on the measurement of the environmental effective amplitude $\delta C_{\rm g}[\%]$ as a function of $C_{\rm g}$, denoted by a red ‘+’ symbol, assuming zero eccentricity and fixed environmental power-law $n_{\rm g}$. The dashed line represents the circular case. The $C_{\rm g}=0$ case is computed using $C_{\rm g}=10^{-18}$ for numerical reasons. We denote the $50$ per cent well-measured threshold by a solid gray line.
@@ 5.1.2 Eccentric MBHB
In the presence of an environment, we compute relative uncertainties for injected signals with eccentricities $e_{0}=0.01$ and $e_{0}=0.1$. In Fig. 2, we show results for $\delta C_{\rm g}[\%]$ and $\delta e_{0}[\%]$ as a function of $C_{\rm g}$. We find that $\delta C_{\rm g}[\%]\approx 6.35\times 10^{3}/|10^{15}C_{\rm g}|$, irrespective of the eccentricity value, as expected, but higher than the circular case shown in Fig. 1. The eccentricity uncertainty, $\delta e_{0}[\%]$, is nearly independent of the environmental amplitude, but reaches higher values than for the vacuum case shown in Table. 2. Both increases can be attributed to having an extra free intrinsic-inspiral parameter. $|C_{\rm g}|\gtrsim 10^{-13}$ is well-measured for our fiducial system.
@@ 5.1.3 Summary
  • For all the scenarios explored here, amplitudes less than $C_{\rm g}]\lesssim 10^{-14}$ are not well measured.
In the next section, we will extend the Fisher results to cover the full parameter space defined in (4). Based on the results in this section, we will only consider $C_{\rm g}\geq 10^{-15}$ and choose $C_{\rm g}=10^{-13}$ and $e_{0}=0.1$ as our fiducial intrinsic-inspiral parameters.
@@ 5.2.1 $\delta C_{\rm g}[\%]$
We want to study the uncertainty in the environmental amplitude $C_{\rm g}$ as a function of $C_{\rm g}$ for both circular and eccentric systems. Based on the findings of Section 5.1, we only need to consider positive values of $C_{\rm g}\in\{10^{-12},10^{-13},10^{-14},10^{-15}\}$ and choose one eccentricity (here $e_{0}=0.1$). In Fig. 3, we show $\delta C_{\rm g}[\%]$ for various different choices of the intrinsic-merger parameters and strength of the environmental effect. Since $M_{z}$ defines if the inspiral part of the signal is in the low or high sensitivity region of the LISA frequency band, it affects the uncertainty on $C_{\rm g}$. Moreover, $q$ and $t_{c}$ set the number of GW cycle and so they also affect the measurement of the environmental amplitude. Based on the parameter dependence in Eq. (18), for $M_{z}=10^{6}~{}{\rm M}_{\sun}$, $q=8.0$ or $\eta=0.1$, and $\xi=100$, we can confidently measure ${\rm f}_{\rm Edd}\gtrsim 0.1$ for circular systems and ${\rm f}_{\rm Edd}\gtrsim 1$ for eccentric MBHBs. These constraints on ${\rm f}_{\rm Edd}$ depend upon setting $\xi=100$, assuming that gas torques become $\mathcal{O}(10)$ stronger near-merger than at the currently explored wider separations where GWs are not important and extrapolating of the simulation results of Dittmann & Ryan (2022), who found increasingly strong torques on binaries for thinner discs (see their Fig. 3).
Figure 3: The relative Fisher errors on the environmental amplitude, $\delta C_{\rm g}[\%]$, as a function of $|C_{\rm g}|$ and the total mass $M_{z}$. The environmental power-law $n_{\rm g}$ is fixed to $4$ in all panels. In all eight panels, we vary $M_{z}$ from $10^{4.5}$ to $10^{6}~{}{\rm M}_{\sun}$ and $C_{\rm g}$ from $10^{-12}$ to $10^{-15}$. For all the left panels, we set $t_{c}=1$ year and assume $t_{c}=4$ years for all the right panels. We have either $q=1.2$ (in the first and third rows) or $q=8.0$ (in the second and fourth rows). In the top four panels, we assume circular orbits and in the bottom four panels we set $e_{0}=0.1$, and allow the initial eccentricity $e_{0}$ to also be a free parameter. We draw a red dashed line and a solid blue line to identify the region of well-measured parameters which have relative errors below $30$ and $50$ per cent, respectively. Moreover, we suppress all errors beyond $100$ per cent. We can at best well measure $C_{\rm g}\gtrsim 10^{-14}$ for circular binaries with $q=8$ and $t_{c}=4$ years and are only able to constrain $C_{\rm g}=10^{-12}$ for an eccentric system with $q=1.2$ and $t_{c}=1$ year.
@@ 5.2.2 $\delta e_{0}[\%]$
Now we want to compute the relative errors on the initial eccentricity, $e_{0}$, as a function of $e_{0}$, with or without the environmental effect. In Fig. 4, we show $\delta e_{0}[\%]$ in vacuum and in the presence of an environment with $C_{\rm g}=10^{-13}$. The measurement on $e_{0}$ strongly depends upon $M_{z}$ for a similar reason to the dependence of $\delta C_{\rm g}[\%]$ upon $M_{z}$. It also varies weakly with $q$ and $t_{c}$, since limited information is contained in the GW cycles observed from the very early inspiral.
Figure 4: The relative error on the initial eccentricity, $\delta e_{0}[\%]$, as a function of $e_{0}$ and $M_{z}$. In all eight panels we vary eccentricities from $10^{-3}$ to $0.1$ and masses from $10^{4.5}$ to $10^{6}~{}{\rm M}_{\sun}$. In the top four panels, we assume vacuum and in the bottom four panels we have $C_{\rm g}=10^{-13}$ and fixed $n_{\rm g}=4$. All other parameter choices are the same as in Fig. 3. In vacuum, we are able to measure at minimum $e_{0}\gtrsim 10^{-2.75}$. While in the presence of an environment, at best we can measure $e_{0}\gtrsim 10^{-2}$.
For $t_{c}=1$ year we find a higher minimum measurable eccentricity in vacuum than what is reported in Garg et al. (2024b) because of the extra number of free intrinsic-merger parameters $\{\chi_{1},\chi_{2},t_{c}\}$ in the analysis.
@@ 6 Bayesian inference
We use the exact same Bayesian setup outlined in Section 4 of Garg et al. (2024b) to find posterior distributions for a few systems of interest. Salient features include a zero noise realization and the Fisher initialization101010starting walkers inside a multivariate Gaussian with its mean given by the injected values and covariance by the Fisher matrix. to accelerate the parallel tempering Markov-chain Monte Carlo (PTMCMC) runs using the ptmcmc [^11] https://github.com/JohnGBaker/ptmcmc sampler. We set uniform priors for all parameters. The resulting posteriors can be used to cross-verify the Fisher matrix results reported in Section 5.
First, we show posteriors for our fiducial system in Fig. 5.121212The posterior for $t_{c}=1$ year is similarly Gaussian as in Fig. 5, although with higher covariances as expected. We can infer that all the parameters are well-recovered and the posteriors almost overlap with the Fisher matrix predictions. We conclude that our Fisher results in Section 5 are robust.  [^13] We also show how posteriors on $C_{\rm g}$ change for different eccentricities in Fig. 9. Moreover, the degeneracy between $e_{0}$ and $C_{\rm g}$ is due to both being inspiral-only effects that decay as the separation decreases. The eccentricity decreases because of the GW-induced circularization and environmental perturbation decays due to entering phase at the $-4$PN order. Over the time that these effects are significant, the orbital velocity does not evolve very much, which allows effects with different post-Newtonian orders to compensate for each other. This means that very similar waveforms are produced for different combinations of $e_{0}$ and $C_{\rm g}$. For instance, increasing environmental amplitude to a higher positive value can be offset by having a higher eccentricity.
@@
Figure 5: Posterior distributions (solid black) for a gas-embedded eccentric MBHB with parameters $M_{z}=10^{5}~{}{\rm M}_{\sun}$, $q=8.0$, $\chi_{1,2}=0.9$, and $t_{c}=4$ years with an effective gas-amplitude $C_{\rm g}=10^{-13}$ and a fixed environmental power-law $n_{\rm g}=4$. The blue lines mark the injected values, the middle dashed line shows the median of the distribution, and the two extreme vertical dashed lines indicate the $90$ per cent symmetric credible interval. The two-dimensional contours of the posteriors indicate $68$, $95$, and $99$ per cent credible intervals. We also show the Fisher matrix predictions (dashed red) for comparison.
Next, in Fig. 6, we show posteriors for the intrinsic-inspiral parameters when the environmental power-law exponent ($n_{\rm g}$) is also a free variable together with all the other intrinsic parameters. We find that keeping $n_{\rm g}$ free leads to non-Gaussian posteriors for $e_{0}$, $C_{\rm g}$, and $n_{\rm g}$, and causes apparent biases in the 1D marginal posteriors for both the power-law exponent and the environmental amplitude. We are using zero noise injections which means that the maximum value of the log-likelihood, i.e., zero, is at the injected parameters. However, this does not mean that there can not be other local maxima in the posterior. There is a degeneracy between $C_{\rm g}$ and $n_{\rm g}$ in the environmental phase correction $\Delta\psi_{\rm gas}$ in Eq. 17. This degeneracy occupies a much larger prior volume around $n_{\rm g}\sim 3.2$ in Fig. 6 than the true peak of the likelihood, which means that even if the likelihood there is lower than at the injected values, the total weight in the evidence could be comparable to that of the true peak. This would mean that the marginal distributions could favour the secondary mode, which is what we are seeing here. Moreover, Fig. 6 implies that a small change in the initial eccentricity can allow a particular value of $n_{\rm g}$ to absorb a large environmental amplitude.
The low SNR in the early inspiral of the signal, where gas dephasing is significant, coupled with the fact that a small variation in the other intrinsic parameters can compensate for relatively larger variations in $C_{\rm g}$ and $n_{\rm g}$, leads to this behavior. We further illustrate this degeneracy in Appendix C. If we bring the source to a redshift $z=0.01$, rather than $z=1$, in order to increase the SNR, then all posteriors become Gaussian as shown in Fig. 10, illustrating that the low SNR in the inspiral phase of the signal is one of the issues. Future high-resolution hydrodynamical simulations of binary-disc interaction with more physics would allow for the identification of richer features and higher-order environmental phase terms, which may help to break this degeneracy at lower SNRs.
Figure 6: Posteriors for $e_{0}$, $C_{\rm g}$, and $n_{\rm g}$, when the environmental power-law is a free parameter (black solid). Posteriors for intrinsic-merger parameters are almost the same as in Fig. 5. For comparison, we also show $e_{0}$ and $C_{\rm g}$ posteriors for our fiducial case where $n_{\rm g}$ is kept fixed (dashed red).
@@
Last, we compare $e_{0}$ and $C_{\rm g}$ posteriors between fixed intrinsic-merger parameters, free intrinsic-merger parameters, and free intrinsic-merger and extrinsic parameters in Fig. 7.141414See Fig. E2 of Garg et al. (2024b) for posteriors of the extrinsic parameters. We find that all posteriors peak around the injected values with smallest uncertainties for the fixed intrinsic-merger variables and largest uncertainties for the free extrinsic variables. These results can be attributed to having fewer or more free parameters. For fixed intrinsic-merger parameters, we have no support for $C_{\rm g}=0$. Also, the errors on the intrinsic parameters are almost independent of the inclusion or not of extrinsic parameters in the model. The case with fixed intrinsic-merger parameters arises when we have independent information on those parameters from either merger-ringdown of the same GW signal or EM counterparts. In reality, we will have narrow priors instead of fixed values for $\{M_{z},q,\chi_{1,2},t_{c}\}$ and posteriors on $e_{0}$ and $C_{\rm g}$ will be somewhere between the black lines and the red dashed lines in Fig. 7.